List the processing steps for the following code.


For decX As Decimal = 6.5 To 8.5
MessageBox.Show(decX.ToString("N1"))
Next decX
Answer #1

Processing Steps:

  • For loop starts with the creation of a counter variable "decX" and assigning the initial value 6.5 to it.
  • It then immediately checks if the value of "decX" is less than or equal to 8.5
    • Since the condition is true, then the value of "decX" is converted to String using the format specifier N1 which displays the decimal number with one decimal value (6.5) in a messagebox as a String.
    • Next decX statement will then increment the value of "decX" by 1, which is the default Step value unless specified otherwise
    • Hence, decX is incremented by 1 and it becomes 7.5
  • The looping construct again check if the value of "decX" is less than or equal to 8.5
    • ​​​​​​​Since the condition is true, the value 7.5 is displayed in Message Box as a string
    • The value of "decX" is incremented by 1
    • "decX" value becomes 8.5
  • ​​​​​​​The looping construct again check if the value of "decX" is less than or equal to 8.5
    • ​​​​​​​Since the the value is now equal to 8.5, the condition results as true and hence the value 8.5 is displayed in Message Box as a string
    • The value of "decX" is incremented by 1
    • "decX" value becomes 9.5
  • ​​​​​​​​​​​​​​The looping construct again check if the value of "decX" is less than or equal to 9.5
    • ​​​​​​​Since the the value is greater than 8.5, the condition results as false and hence for loop stops executing the statements in its body
    • The flow of program control will move on to the next statement following the for loop.
    • In other words, the program continues with statements out of for loop
Current value of decX if decX is less than or equal to 8.5 Statement in for loop Next value of decX
Initially, decX = 6.5 if 6.5 <= 8.5 , TRUE displays 6.5 in messagebox 6.5 + 1 = 7.5
decX = 7.5 if 7.5 <= 8.5 , TRUE displays 7.5 in messagebox 7.5 + 1 = 8.5
decX = 8.5 if 8.5 <= 8.5 , TRUE displays 8.5 in messagebox 8.5 + 1 = 9.5
decX = 9.5 if 7.5 <= 8.5 , FALSE LOOP STOPS